# TOPA: Extend Large Language Models for Video Understanding via Text-Only Pre-Alignment

### v1

@@ Abstract
Recent advancements in image understanding have benefited from the extensive use of web image-text pairs. However, video understanding remains a challenge despite the availability of substantial web video-text data. This difficulty primarily arises from the inherent complexity of videos and the inefficient language supervision in recent web-collected video-text datasets. In this paper, we introduce Text-Only Pre-Alignment (TOPA), a novel approach to extend large language models (LLMs) for video understanding, without the need for pre-training on real video data. Specifically, we first employ an advanced LLM to automatically generate Textual Videos comprising continuous textual frames, along with corresponding annotations to simulate real video-text data. Then, these annotated textual videos are used to pre-align a language-only LLM with the video modality. To bridge the gap between textual and real videos, we employ the CLIP model as the feature extractor to align image and text modalities. During text-only pre-alignment, the continuous textual frames, encoded as a sequence of CLIP text features, are analogous to continuous CLIP image features, thus aligning the LLM with real video representation. Extensive experiments, including zero-shot evaluation and finetuning on various video understanding tasks, demonstrate that TOPA is an effective and efficient framework for aligning video content with LLMs. In particular, without training on any video data, the TOPA-Llama2-13B model achieves a Top-1 accuracy of 51.0% on the challenging long-form video understanding benchmark, Egoschema. This performance surpasses previous video-text pre-training approaches and proves competitive with recent GPT-3.5-based video agents.
@@ 1 Introduction
Image-language understanding has made large advancements in both image-language alignment [22, 43] and Multimodal Large Language Models (MLLMs) [1, 23, 29, 79], benefiting from pre-training on large-scale noise-paired image-text data collected from the web [6, 17, 45, 47, 46]. This raises a question: Can we mirror this success in video-language understanding? Research [40, 56, 67, 76] has explored pretraining video-language models on millions of web video-text data [3, 35], achieving promising results in basic video tasks such as video-text retrieval, video captioning, and video question answering across conventional video benchmarks. However, recent research reveals that these models struggle with the challenging long-form video understanding benchmark, i.e., Egoschema [34], which requires intrinsic temporal understanding capabilities. This highlights the gap in adapting web video-text pretrained models to more comprehensive video understanding tasks.
@@
We attribute this gap to two primary factors: 1) The intrinsic complexity of the video modality. Videos introduce intrinsic complexities in both spatial and temporal dimensions, which are absent in static images. These complexities require extensive training on larger-scale data to effectively capture video dynamics. Furthermore, representing videos typically involves processing multiple frames, significantly increasing computational demands compared to image modeling. The dual challenges of large-scale training and increased computational requirements make video-language modeling particularly challenging. 2) The limitations of web language supervision. The language supervision in recent web video-text datasets primarily comes from subtitles associated with the videos. These subtitles are often frame-level descriptions that lack the necessary temporal associations crucial for video understanding. Moreover, this form of descriptive supervision is inefficient in building robust reasoning capabilities, especially in terms of temporal reasoning. This mismatch between the richness of video content and the limited language supervision hinders effective video-language modeling.
@@
In this paper, we propose an innovative approach to develop video understanding capabilities using LLMs to simulate video dynamics. Instead of directly aligning video with language, we introduce a textual video representation — a sequence of textual frames designed to mimic real visual dynamics. This textual video can be readily generated by advanced LLMs and effectively simulates various video dynamics by describing them in text. Specifically, we present a Textual Video (TextVid) dataset, automatically generated by LLMs. TextVid includes: 1) Textual videos (hereinafter referred to as “Tideo”), which consist of a sequence of textual frames crafted to mimic the keyframes of real videos, and 2) Tideo annotations, including comprehensive Tideo-level dense descriptions and varied question-answer (QA) pairs. These annotations are of high quality and closely align with the Tideo content, by virtue of the powerful capability of LLM in language generation.
@@
Building on the proposed TextVid dataset, we introduce the Text-Only Pre-Alignment (TOPA) framework, to effectively and efficiently pre-align LLMs with the video modality, reducing the need for costly video-text pre-training. We introduce three tasks for video-LLM pre-alignment: Tideo summarization, Tideo QA and multi-choice Tideo QA. To bridge the gap between textual Tideos and visual videos, we leverage the CLIP model for feature extraction. Specifically, we employ the CLIP text encoder to extract frame-level representations for Tideos, and the CLIP visual encoder for real videos. During the text-only pre-alignment phase, the LLM learns to process continuous CLIP text features. In the real video inference phase, it transitions to handling continuous CLIP image features. Due to the aligned CLIP image-text feature space, the LLM can adapt to real video inputs despite being trained on textual representations. Our main contributions include:
@@
(3) Extensive experiments demonstrate TOPA’s effectiveness across various video understanding tasks. Particularly, the TOPA-Llama2-13B model achieves 51.0% Top 1 accuracy in the challenging EgoSchema benchmark, outperforming previous video-text pretraining methods and approaching recent GPT-4-based video agents.
@@ 2 Related Work
Vision-Language Alignment. Radford et al. [43] align the vision and language in a common feature space via contrastive learning with large-scale web image-text data. [1, 23, 29, 79] align the visual model with advanced LLM via training on image-caption pairs and interleaved image-text data. [7, 19, 28, 73] explore modeling video sequences within LLM spaces, leveraging LLM for video-language understanding. In this paper, we focus on video-LLM alignment. Rather than using multimodal data for vision-language alignment, we introduce a novel text-only pre-alignment framework to extend LLMs for video understanding without pre-training on real video-text data.
LLMs for multimodal data augmentation. Recent research explores to leverage LLMs to enhance the multimodal data. A line of work [5, 12, 29] use LLMs for refining captions or extending the image caption pairs to diverse visual tasks like visual conversation and image editing. Another line of work [24, 32, 33, 42] further employ advanced LLM to enrich web video supervision for video instruction tuning. In this paper, rather than enhancing multimodal datasets, we propose generating text-only data consisting of "textual videos" and diverse language supervision, which aims to simulate real videos and their corresponding annotations.
Long-form Video Understanding. Long-form video understanding presents significant challenges due to the intricate spatial and temporal dynamics. Conventional video-text pretraining approaches [4, 40, 57, 58, 67, 80] utilize extensive web video-caption data for video-language alignment. Another line of research [18, 48, 71] seeks to adapt recent image MLLMs to video understanding. Recent research [24, 58, 73, 78] employ video instruction-tuning for video-LLM alignment to enhance video-language understanding. Recently, [9, 37, 49, 53, 55, 68, 72] combine the LLM with various VLM tools as video agents to perform video-understanding tasks. In this paper, we propose a novel text-only pre-alignment framework to efficiently and effectively align LLMs with videos without pre-training on real videos.
@@ 3.1 TextVid Dataset
This dataset, comprising textual videos (Tideos) and associated annotations, is generated by an advanced LLM (i.e., Gemini Pro 1.0 [50]). The data generation pipeline is detailed in Appendix B. Each Tideo is presented in a textual format and contains 5-15 sequential frames. Each frame includes a frame caption that describes the scene and multiple object captions. To enhance understanding and interaction with these Tideos, the dataset features a dense description summarizing the Tideo, as well as a set of multiple-choice questions and answers related to the Tideo content. The structure of each element is as follows:
@@
There are two major advantages of the TextVid dataset. (1) The large-scale and diverse Tideos. As the dataset is text-only and fully generated by an LLM, the size of TextVid is unlimited. Moreover, the Tideos can cover a broad range of domains by simply prompting the language model with appropriate conditions. It is distinctly different from previous web video-text dataset like Howto100M [35] that are limited to specific human-centric instructional videos. In practice, we enhance the diversity of TextVid by randomly sampling video captions from WebVid-2M [3], video titles from Howto100m [35], video tasks from Ego4D [14] and object names with descriptions from WordNet [36] as a condition of prompts. These varied prompts enable the language model to generate a diverse dataset. (2) The high-quality, consistent and free-form language supervision. The language supervisions are generated along with Tideos. The advanced capabilities of LLM ensure the quality of these supervisions, making them less noisy than web video-text data. Moreover, both the Tideo and the supervision are in textual format, making the supervision closely aligned with the Tideo’s content. Additionally, the format of the language supervision is unrestricted. For example, we prompt an LLM to generate dense description and multi-choice QA pairs as language supervision in TextVid.
@@ 3.2 Text-only Pre-alignment
Video-LLM alignment. The goal of video-LLM alignment is to extend the language-only LLMs for processing video inputs. Given a video sampled with $n$ frames $\{\mathbf{I}_{1},\mathbf{I}_{2},\ldots,\mathbf{I}_{n}\}$, we follow recent work [19, 67] to leverage a frozen CLIP model to extract the frame-level visual feature, formulated as $\mathbf{f}^{v}_{i}={E}_{\text{image}}(\mathbf{I}_{i})$, where $E_{\text{image}}$ denotes CLIP image encoder. The CLIP features are then projected into the LLM space via a simple linear layer, denoted as ${G}({P}(\mathbf{f}^{v}_{1}),...,{P}(\mathbf{f}^{v}_{n}))$, where ${G}$ denotes a language model and ${P}$ denotes a projection layer that projects the CLIP feature to LLM space.
@@
Tideo Representation. In this work, we leverage Tideos (c.f. Section 3.1) for video-LLM pre-alignment instead of training on real videos. Specifically, given the textual frame $T_{i}$, we employ CLIP text encoder to extract the frame representation from frame caption $C_{i}$ and detailed object captions $D_{i}$, represented as $\mathbf{f}^{t}_{i}={F}_{\text{fusion}}({E}_{\text{text}}(C_{i}),{E}_{\text{text}}(D_{i,1}),...,{E}_{\text{text}}(D_{i,j}))$, where $F_{\text{fusion}}$ is a fusion function such as simple average pooling, and $E_{\text{text}}$ denotes the CLIP text encoder. A Tideo with $n$ textual frames is represented as $\mathbf{V}^{t}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{n}\}$
@@
Text-only Training. Given the Tideo $T$, dense Tideo-level description $D_{V}$, and QA pairs with multiple choices $\{(Q_{k},O_{k},A_{k})\}$, we introduce the following tasks for Tideo-LLM alignment: (1) Tideo Summarization: Given the Tideo, generate a detailed description to summarize the Tideo; (2) Tideo QA: Given the Tideo and question, predict the answer; (3) Multi-choice Tideo QA: Given the Tideo, question and multiple choices, choose the correct answer from the candidates. We employ a unified auto-regressive Language Modeling (LM) objective for these three tasks:
$\mathcal{L}_{\text{LM}}(\theta_{{G}},\theta_{P})=-\frac{1}{|t|}\sum_{i=1}^{|t|}\log{G}\big{(}t_{i}|{P}(\mathbf{V}),{Z},t_{<i}\big{)},$ (1)
@@
where $\mathbf{V}$ denotes the Tideo representation, and $\mathbf{V}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{i}\}$ during the text-only training, Z denotes the task specific condition tokens and $t_{i}$ denotes the $i_{th}$ target token. $\theta_{{G}}$ and $\theta_{{P}}$ denote the learnable parameters of the LLM adapter and the projection layer $P$, respectively. In practice, we use the following format as the LLM input: {Task Instruction}. Video:$\mathbf{f}^{t}_{1}$, …, $\mathbf{f}^{t}_{i}$}. {Task Conditions}. Answer: {Predict Targets}. For the Tideo summarization task, the target is detailed Tideo descriptions. For Tideo QA task, the target is the answer and the condition is the question. For multi-choice Tideo QA task, the target is the correct option and the condition consists with question and options. The details of the task-specific prompts are included in Appendix D.1.
3.3 Zero-shot Inference
@@
Section 3.2 introduces the text-only pre-alignment using the TextVid dataset. In this section, we detail how to adapt this text-only pre-aligned LLM for real videos. During pre-alignment, we leverage the textual representation $\mathbf{V}^{t}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{i}\}$ as the Tideo representation. During inference, we take real videos features as input, i.e., $\mathbf{V}^{v}=\{\mathbf{f}^{v}_{1},...,\mathbf{f}^{v}_{i}\}$, where $\mathbf{f}^{v}_{i}=E_{\text{image}}(\mathbf{I}_{i})$. These two modality features $\mathbf{f}^{t}$ and $\mathbf{f}^{v}$ that come from CLIP image encoder and CLIP text encoder are aligned via CLIP pre-training. This aligned image-text representation makes it possible to perform zero-shot inference without additional finetuning. However, the modality gap phenomenon [15, 26, 27, 39, 75], i.e., CLIP image feature and CLIP text feature are located in two completely separate regions of the feature space, prevents us from directly taking the visual feature $\mathbf{f}^{v}$ as the textual feature $\mathbf{f}^{t}$. To bridge this modality gap, we follow DeCap [26] to employ a support memory to project the CLIP visual feature into the CLIP text feature space. This training-free projection process is formulated as:
@@
where $\mathbf{m}_{i}$ denotes CLIP text feature from a pre-constructed memory of size $N$, $\mathbf{f}^{v}$ denotes input frame feature of real video and $\mathbf{f}^{v\rightarrow t}$ denotes the projected feature. During zero-shot inference, we take the $\mathbf{V}^{{v\rightarrow t}}=\{\mathbf{f}^{v\rightarrow t}_{1},...,\mathbf{f}^{{v\rightarrow t}}_{i}\}$ as the real video’s representation.

### v2

# TOPA: Extending Large Language Models for Video Understanding via Text-Only Pre-Alignment

@@ Abstract
Recent advancements in image understanding have benefited from the extensive use of web image-text pairs. However, video understanding remains a challenge despite the availability of substantial web video-text data. This difficulty primarily arises from the inherent complexity of videos and the inefficient language supervision in recent web-collected video-text datasets. In this paper, we introduce Text-Only Pre-Alignment (TOPA), a novel approach to extend large language models (LLMs) for video understanding, without the need for pre-training on real video data. Specifically, we first employ an advanced LLM to automatically generate Textual Videos comprising continuous textual frames, along with corresponding annotations to simulate real video-text pairs. Then, these annotated textual videos are used to pre-align language-only LLMs with the video modality. To bridge the gap between textual and real videos, we employ the CLIP model as the feature extractor to align image and text modalities. During text-only pre-alignment, the continuous textual frames, encoded as a sequence of CLIP text features, are analogous to continuous CLIP image features, thus aligning the LLM with real video representation. Extensive experiments, including zero-shot evaluation and finetuning on various video understanding tasks, demonstrate that TOPA is an effective and efficient framework for aligning video modality with LLMs. In particular, without training on any video data, the TOPA-Llama2-13B model achieves a Top-1 accuracy of 51.0% on the challenging long-form video understanding benchmark, EgoSchema. This performance surpasses previous video-text pre-training approaches and is competitive with recent GPT-3.5-based video agents.
@@ 1 Introduction
Image-language understanding has made large advancements in both image-language alignment [26, 48] and Multimodal Large Language Models (MLLMs) [1, 27, 34, 89], aided by pre-training on large-scale noise-paired image-text data collected from the web [6, 19, 51, 53, 52]. This raises a question: Can we mirror this success in video-language understanding? Research [45, 64, 77, 86] has explored pretraining video-language models on millions of web video-text data [3, 40, 65], achieving promising results in basic video tasks such as video-text retrieval, video captioning, and video question answering across conventional video benchmarks. However, recent research reveals that these models struggle with a challenging long-form video understanding benchmark, i.e., EgoSchema [39], which requires intrinsic temporal understanding capabilities. This highlights the gap in adapting web video-text pretrained models to more comprehensive video understanding tasks.
@@
We attribute this gap to two primary factors: 1) The intrinsic complexity of the video modality. Videos introduce intrinsic complexities in both spatial and temporal dimensions, which are not present in static images. These complexities require extensive training on larger-scale data to effectively capture video dynamics. Furthermore, representing videos typically involves processing multiple frames, significantly increasing computational demands compared to image modeling. The dual challenges of large-scale training and increased computational requirements make video-language modeling particularly challenging. 2) The limitations of web language supervision. The language supervision in recent web video-text datasets primarily comes from subtitles or descriptions associated with the videos [3, 40]. However, subtitles often suffer from the issues of visual-textual misalignment [33, 17]. Moreover, the form of descriptive supervision is inefficient in building robust video reasoning capabilities, especially in terms of temporal reasoning. This mismatch between the complex video content and the limited supervision hinders effective video-language modeling.
@@
In this paper, we propose an innovative approach to develop video understanding capabilities by using LLMs to simulate and understand video dynamics. Instead of directly aligning LLMs with real video representation, we first introduce a textual video representation — a sequence of textual frames designed to mimic real visual dynamics. This textual video can be readily generated by advanced LLMs and effectively simulates various video dynamics by describing them in text. Specifically, we present a Textual Video (TextVid) dataset, automatically generated by LLMs. TextVid includes: 1) Textual videos (hereinafter referred to as “Tideo”), which consist of a sequence of textual frames crafted to mimic the keyframes of real videos, and 2) Tideo annotations, including comprehensive Tideo-level dense descriptions and varied question-answer (QA) pairs. These annotations are of high quality and closely align with the Tideo content, by virtue of the powerful capability of LLM in language generation.
@@
Building on the proposed TextVid dataset, we introduce the Text-Only Pre-Alignment (TOPA) framework, to effectively and efficiently pre-align LLMs with the video modality, reducing the need for costly video-text pre-training. We introduce three tasks for video-LLM pre-alignment: Tideo summarization, Tideo QA and multi-choice Tideo QA. To bridge the gap between textual Tideos and visual videos, we leverage the CLIP [48] model for feature extraction. Specifically, we employ the CLIP text encoder to extract frame-level representations for Tideos, and the CLIP visual encoder for real videos. During the text-only pre-alignment phase, the LLM learns to process continuous CLIP text features of Tideos. In the real video inference phase, it transitions to handling continuous CLIP image features of real video. Due to the aligned CLIP image-text feature space, the LLM can adapt to real video inputs despite being trained on textual representations. Our main contributions include:
@@
(3) Extensive experiments demonstrate TOPA’s effectiveness across various video understanding tasks. Particularly, the TOPA-Llama2-13B model achieves 51.0% Top-1 accuracy in the challenging EgoSchema benchmark, outperforming previous video-text pretraining methods and competitive with recent GPT-3.5-based video agents.
@@ 2 Related Work
Vision-language alignment. CLIP [48] aligns the vision and language modalities in a common feature space via contrastive learning with large-scale web image-text data. MLLMs [1, 27, 34, 89] align the visual model with LLM via training on image-caption pairs and interleaved image-text data. Video-LLMs [7, 23, 32, 83] explore modeling video sequences within LLM spaces, leveraging LLM for video-language understanding. In this paper, we focus on video-LLM alignment. Rather than using multimodal data for vision-language alignment, we introduce a novel text-only pre-alignment framework to extend LLMs for video understanding without pre-training on real video-text data.
LLMs for multimodal data augmentation. Recent research explores the use of LLMs to enhance the multimodal data. A line of work [5, 12, 34] use LLMs for refining captions or extending the image caption pairs to diverse visual tasks like visual conversation and image editing. Another line of work [28, 37, 38, 47] further employ advanced LLM to enrich web video supervision for video instruction tuning. In this paper, rather than enhancing multimodal datasets, we propose generating text-only data consisting of "textual videos" and diverse language supervision, which aims to simulate real videos and their corresponding annotations.
Long-form video understanding. Long-form video understanding [39, 61, 71] presents significant challenges due to the intricate spatial and temporal dynamics. Conventional video-text pretraining approaches [4, 45, 65, 66, 77, 90] utilize extensive web video-caption data for video-language alignment. Recent research [28, 66, 83, 88] employ video instruction-tuning for video-LLM alignment to enhance video-language understanding. Another line of research [22, 54, 81, 49] seeks to adapt recent image MLLMs to video understanding. A parallel line of research [9, 42, 55, 60, 63, 13, 78, 82, 67, 21] combine the LLM with various VLM tools as video agents to perform video-understanding tasks. In this paper, we propose a novel text-only pre-alignment framework to efficiently and effectively align LLMs with videos without pre-training on real videos.
@@ 3.1 TextVid Dataset
This dataset, comprising textual videos (Tideos) and associated annotations, is generated by an advanced LLM (i.e., Gemini Pro 1.0 [56]). The data generation pipeline is detailed in Appendix D. Each Tideo is presented in a textual format and contains 5-15 sequential frames. Each frame includes a frame caption that describes the scene and multiple object captions. To enhance understanding and interaction with these Tideos, the dataset features a dense description summarizing the Tideo, as well as a set of multiple-choice questions and answers related to the Tideo content. The structure of each element is as follows:
@@
There are two major advantages of the TextVid dataset. (1) The large-scale and diverse Tideos. As the dataset is text-only and fully generated by an LLM, the size of TextVid is practically unlimited. Moreover, the Tideos can cover a broad range of domains by simply prompting the language model with appropriate conditions. It is distinctly different from previous web video-text dataset like Howto100M [40] that are limited to specific human-centric instructional videos. In practice, we enhance the diversity of TextVid by randomly sampling video captions from WebVid-2M [3], video titles from Howto100m [40], video tasks from Ego4D [15] and object names with descriptions from WordNet [41] as a condition of prompts. These varied prompts enable the language model to generate a diverse dataset. (2) The high-quality, consistent and free-form language supervision. The language supervisions are generated along with Tideos. The advanced capabilities of LLM ensure the quality of these supervisions, making them less noisy than web video-text data. Moreover, both the Tideo and the supervision are in textual format, making the supervision closely aligned with the Tideo’s content. Additionally, the format of the language supervision is unrestricted. For example, we prompt the LLM to generate dense descriptions and multi-choice QA pairs as language supervision.
@@ 3.2 Text-Only Pre-Alignment
Preliminary: Video-LLM alignment. The goal of video-LLM alignment is to extend pre-trained LLMs for processing video inputs. Given a video sampled with $n$ frames $\{\mathbf{I}_{1},\mathbf{I}_{2},\ldots,\mathbf{I}_{n}\}$, Recent work [23, 77] uses a frozen CLIP model to extract the frame-level visual feature, formulated as $\mathbf{f}^{v}_{i}={E}_{\text{image}}(\mathbf{I}_{i})$, where $E_{\text{image}}$ denotes CLIP image encoder. The CLIP features are then projected into the LLM space via a simple linear layer, denoted as ${G}({P}(\mathbf{f}^{v}_{1}),...,{P}(\mathbf{f}^{v}_{n}))$, where ${G}$ denotes a language model and ${P}$ denotes a projection layer that projects the CLIP feature to LLM space.
@@
Tideo representation. In this work, we leverage Tideos (c.f. Section 3.1) for video-LLM pre-alignment instead of training on real videos. Specifically, given the textual frame $T_{i}$, we employ CLIP text encoder to extract the frame representation from frame caption $C_{i}$ and detailed object captions $D_{i}$, represented as $\mathbf{f}^{t}_{i}={F}_{\text{fusion}}({E}_{\text{text}}(C_{i}),{E}_{\text{text}}(D_{i,1}),...,{E}_{\text{text}}(D_{i,j}))$, where $F_{\text{fusion}}$ is a fusion function such as simple average pooling, and $E_{\text{text}}$ denotes the CLIP text encoder. A Tideo with $n$ textual frames is represented as $\mathbf{V}^{t}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{n}\}$.
@@
Text-only pre-alignment. Given the Tideo $T$, dense Tideo-level description $D_{V}$, and QA pairs with multiple choices $\{(Q_{k},O_{k},A_{k})\}$, we introduce the following tasks for Tideo-LLM alignment: (1) Tideo Summarization: Given the Tideo, generate a detailed description to summarize the Tideo; (2) Tideo QA: Given the Tideo and question, predict the answer; (3) Multi-choice Tideo QA: Given the Tideo, question and multiple choices, choose the correct answer from the candidates. We employ a unified auto-regressive Language Modeling (LM) objective for these three tasks:
$\mathcal{L}_{\text{LM}}(\theta_{{G}},\theta_{P})=-\frac{1}{|t|}\sum_{i=1}^{|t|}\log{G}\big{(}t_{i}|{P}(\mathbf{V^{t}}),{Z},t_{<i}\big{)},$ (1)
@@
where $\mathbf{V^{t}}$ denotes the Tideo representation, and $\mathbf{V^{t}}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{n}\}$ during the text-only training, Z denotes the task specific condition tokens and $t_{i}$ denotes the $i_{th}$ target token. $\theta_{{G}}$ and $\theta_{{P}}$ denote the learnable parameters of the LLM adapter and the projection layer $P$, respectively. In practice, we use the following format as the LLM input: {Task Instruction}. Video:$\mathbf{f}^{t}_{1}$, …, $\mathbf{f}^{t}_{n}$}. {Task Conditions}. Answer: {Predict Targets}. For the Tideo summarization task, the target is detailed Tideo descriptions. For Tideo QA task, the target is the answer and the condition is the question. For multi-choice Tideo QA task, the target is the correct option and the condition consists with question and options. The details of the task-specific prompts are included in Appendix F.1.
3.3 Adapting to Real Video Understanding
Section 3.2 introduces the text-only pre-alignment using the TextVid dataset. In this section, we detail how to adapt this text-only pre-aligned LLM for real video understanding. We introduce two approaches: one is zero-shot inference, which directly infers with real video data. And the other is supervised finetuning, where the pre-aligned model is further finetuned on downstream video data.
@@ 3.3 Adapting to Real Video Understanding
Zero-shot inference. During pre-alignment, we leverage the textual representation $\mathbf{V}^{t}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{n}\}$ as the Tideo representation. During inference, we take real videos features as input, i.e., $\mathbf{V}^{v}=\{\mathbf{f}^{v}_{1},...,\mathbf{f}^{v}_{n}\}$, where $\mathbf{f}^{v}_{i}=E_{\text{image}}(\mathbf{I}_{i})$. These two modality features $\mathbf{f}^{t}$ and $\mathbf{f}^{v}$ that come from CLIP image encoder and CLIP text encoder are aligned via CLIP pre-training. This aligned image-text representation makes it possible to perform zero-shot inference without additional finetuning. However, the modality gap phenomenon [16, 30, 31, 44, 85], i.e., CLIP image feature and CLIP text feature are located in two completely separate regions of the feature space, prevents us from directly taking the visual feature $\mathbf{f}^{v}$ as the textual feature $\mathbf{f}^{t}$. To bridge this modality gap, we follow DeCap [30] to employ a support memory to project the CLIP visual feature into the CLIP text feature space. This training-free projection process is formulated as:
@@
where $\mathbf{m}_{i}$ denotes CLIP text feature from a pre-constructed memory of size $N$, $\mathbf{f}^{v}$ denotes input frame feature of real video and $\mathbf{f}^{v\rightarrow t}$ denotes the projected feature. During zero-shot inference, we take the $\mathbf{V}^{{v\rightarrow t}}=\{\mathbf{f}^{v\rightarrow t}_{1},...,\mathbf{f}^{{v\rightarrow t}}_{n}\}$ as the real video’s representation.
Supervised finetuning. On the other hand, the text-only pre-alignment can be viewed as a pretraining stage. Following the pretraining-finetuning paradigm, the pre-aligned LLMs can then be fine-tuned on real video data for improved downstream task performance. The finetuning process is similar to the text-only pre-alignment as detailed in Section 3.2, except that the LLM receives a sequence of CLIP visual features as input instead of CLIP textual features.